40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Jan 14 00:50:08 2012, rana, koji, Configuration, IOO, Towards coating thermal noise measurement with RefCav / MC beat 
    Reply  Tue Jan 17 03:10:27 2012, kiwamu, Configuration, IOO, rotated lambda/2 plate  
Message ID: 6198     Entry time: Sat Jan 14 00:50:08 2012     Reply to this: 6205
Author: rana, koji 
Type: Configuration 
Category: IOO 
Subject: Towards coating thermal noise measurement with RefCav / MC beat 

Koji asked aloud tonight if we could measure the coating thermal noise of the refcav optics by beating the refcav light with the MC_TRANS light. Then we looked at our calculations for the noises:

Displacement noise of T=200ppm silica/tantala coating on a 1" silica substrate with a 300 micron beam spot = 1e-18 * sqrt(100 Hz / f) m/rHz.

Displacement noise from coating thermal in the MC is roughly smaller by the beam size ratio (1.8 mm / 0.3 mm). Some differences due to 3 mirrors and more layers on MC2 than the others, but those are small factors.

So, the frequency noise from the refcav should by larger than the MC thermal noise by a total factor of (1.8 / 0.3) * (13 m / 8 inches) ~ 400.

Another way to say it is that the effective strain noise in the RC is (1e-18 / 0.200) = 5e-18 /rHz. This translates into (5e-18 * 13) = 6.5e-17 m/rHz in the MC. (in frequency noise its 1.5 mHz/rHz).

I have measured the frequency noise in the LLO MC to be at this level back in 2009, so it seems possible to use our RC + MC to measure coating thermal noise by the length amplification factor and compete with Frank+Tara.

 

So today we set up the Jenny RC temperature setup to lock the LWE NPRO to the RC and then set up the beat note with the IFO REFL beam on the AS table. By using the 2 laser beat, we are avoiding the VCO phase noise issue which used to limit the PSL frequency noise at ~0.01 Hz/rHz. To do this we have reworked some of the optics on the PSL and AS tables, but I think its been done without disturbing the beams for the regular locking. Beat note has been found, but the NPRO has still not been locked to the RC - next we setup the lockin amp, dither the PZT, and then use the New Focus lock box to lock it to the RC.

You might think that its hard to measure this since the MC has ~1 MHz frequency fluctuations and we want to measure down to 1e-4 Hz. But, in fact, we can just use a 200 m MFD with a LT1128 preamp. Then we use the MFD to stabilize the MC length to the refcav and just use the control + error signal of the MFD setup as the coating thermal noise measurement.

 

Note: Beat found at ~40deg for the aux laser. The aux laser is on but the shutter is closed.
The AS camera seems to be hosed. Need a bit of alignment. (KA) ==> Fixed. (Jan 15)

ELOG V3.1.3-